OpenStudy (anonymous):

Simplify the expression 2/x-6+4x

jigglypuff314 (jigglypuff314):

do you mean \[\frac{ 2 }{ x }-6+4x\] or \[\frac{ 2 }{ x-6 }+4x\] or \[\frac{ 2 }{ x-6+4x }\]

OpenStudy (anonymous):

I mean the 2nd equation you listed

jigglypuff314 (jigglypuff314):

i supposed you would give them a common denominator so \[\frac{ 2 }{ x-6 }+4x \rightarrow \frac{ 2 }{ x-6 }+\frac{ 4x(x-6) }{ x-6 }\] then add and distribute so \[\frac{ 2+4x(x-6) }{ x-6 } \rightarrow \frac{ 2+4x ^{2}-24x }{ x-6 } \rightarrow \frac{ 4x ^{2}-24x+2 }{ x-6 }\]

OpenStudy (anonymous):

Oh ok that makes more sense then the answer I had gotten. Would you mind helping me with another one? I am not positive that I got the right answer.

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

ok thank you! it is x/x+5+5x/x-5 It is the same kind of one that you did above.

jigglypuff314 (jigglypuff314):

you mean \[\frac{ x }{ x+5 } + \frac{ 5x }{ x-5 }\]

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

multiply each value by\[\frac{ (x+5)(x-5) }{ (x+5)(x-5)}\]

jigglypuff314 (jigglypuff314):

so you get \[\frac{ x(x-5) }{ (x+5)(x-5) }+ \frac{ 5x(x+5) }{ (x+5)(x-5) }\] then distribute and add like the previous problem

jigglypuff314 (jigglypuff314):

my mistake, correction: multiply each value by (x+5)/(x+5) or (x-5)/(x-5) as needed to give the values a common denominator so that you can add them together.

OpenStudy (anonymous):

i got 6-5x?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

do you know how to distribute?

OpenStudy (anonymous):

I think i am distributing wrong, could you show me please?

jigglypuff314 (jigglypuff314):

\[x(x-5) \rightarrow x ^{2}-5x\]

jigglypuff314 (jigglypuff314):

\[5x(x+5) \rightarrow 5x ^{2}+25x\]

OpenStudy (anonymous):

ok so now I got 6x^2+20x/x^2-25?

jigglypuff314 (jigglypuff314):

yes

OpenStudy (anonymous):

do I need to simplify it even more?

jigglypuff314 (jigglypuff314):

I am not sure what sort of math class you are in nor do I know what your teacher expects but I don't think you can simplify it even more.

OpenStudy (anonymous):

Ok thank you very much for your help!

jigglypuff314 (jigglypuff314):

your welcome :)